OpenStudy (anonymous):

Which expression is equivalent to tan(-x)cos2x? A. -sin xcos x B. sin xcos x C. tan x + sin2x D. tan x − sin2x E. -tan x − sin2x

OpenStudy (anonymous):

@souvik i dont understand

OpenStudy (souvik):

do you know this ...tan(-x)=-tanx?

OpenStudy (anonymous):

yea

OpenStudy (souvik):

and this cos2x=cos^2x-sin^2x...........?

OpenStudy (anonymous):

yes

OpenStudy (souvik):

then use these formulas...

OpenStudy (anonymous):

i dont know how to exactly

OpenStudy (souvik):

tan(-x)*cos2x =-tanx[cos^2x-sin^2x]

OpenStudy (anonymous):

i got that but then what?

OpenStudy (souvik):

then tanx=sinx/cosx sinx/cosx[cos^2x-sin^2x]

OpenStudy (anonymous):

what now do they cancel out or something?

OpenStudy (souvik):

\[-\frac{ sinx }{ cosx }*\cos^2x +\frac{ sinx }{ cosx }*\sin^2x\]

OpenStudy (anonymous):

-sinxcosx+tan(x)? thats not one of the choices

OpenStudy (anonymous):

@souvik?

OpenStudy (anonymous):

you there??

OpenStudy (souvik):

yeah i am here..

OpenStudy (anonymous):

-sinxcosx+tan(x) is what i got

OpenStudy (souvik):

we should do it in this way... using this...cos2x=2cos^2x-1

OpenStudy (souvik):

-tanx[2cos^2x-1] =-sinx/cosx[2cos^2x-1] =-2sinx*cosx+sinx/cosx =-sin2x+tanx =tanx-sin2x get it?

OpenStudy (anonymous):

kinda, buti think i will get better with more practice problems. thank you for your help.

OpenStudy (souvik):

the more u practice the more u learn......:)